# Perceptually Optimized Super Resolution

### v1

@@ 1 Introduction
Super-resolution (SR) has quickly become a fundamental tool in imaging and media distribution, given the increasing requirements of delivering higher quality content at lower bandwidths, and as general compression tools to deal with escalating imaging sensor resolutions. In media production such as virtual reality, augmented reality, and video games, SR is indispensable to cater to the high-resolution, high-framerate requirements of modern displays and low power budgets [32, 30, 2, 15]. With the advent of hardware-specific accelerators for efficiently running DL models [31], most modern approaches, even for real-time needs, involve using convolutional neural networks trained on large priors of natural images, which during run-time take low-resolution images as input and produce higher resolution versions. However, even with ever-increasing computational power at our disposal, the computational burden of high-quality SR methods is still problematic [1]. In fact, state-of-the-art methods on real-time SR using neural networks, such as Bicubic++ [7], directly compete with efficient, classic bicubic interpolation techniques, yielding marginal quantitative improvements while still struggling to compete in runtime. Our key insight is that, given that SR images are to be observed by a human, we can take advantage of the naturally compressive capabilities of the human visual system (HVS) to process differently areas of the input image depending on its characteristics and human sensitivity to those, such as spatial frequency, luminance, color, contrast, or even, in the case of videos, motion. Computational resources are allocated to perceptually meaningful areas, as determined by our low-level visual model; analogously to how lossy compression schemes such as JPEG similarly allocate memory by leveraging the HVS.
@@
This observation is the foundation of our work. First, we model the output attenuation (in the frequency domain) of baseline SR neural networks after each layer. Then, after dividing an image or individual frame into a set of patches, our perceptual model (grounded on recent work on contrast-sensitivity functions [28]) predicts how many layers of the network should the patch be processed through; perceptually sensitive patches will be processed by the whole model, whereas the least perceptually meaningful content will simply resort to bicubic interpolation, Fig. 1 shows a visual example. We apply our model to other use cases (with the goal of computational efficiency) such as selecting a network of appropriate depth from a set of candidates. In all of these cases, the goal is make the inference as fast as possible without noticeable quality degradation. Through a series of user studies and quantitative measurements of quality and runtime, we demonstrate that our perceptual model enables faster runtime at no perceivable loss of quality when compared with the baseline models. Our work can be integrated into any learned super-resolution architecture, and due to its eccentricity-aware vision model, can be particularly relevant to VR and AR applications, where computational efforts are further directed primarily towards areas in the foveal vision, the foundation of foveated rendering approaches [32].
Figure 1: Visual results of our method compared to the original networks on images 829 and 843 from the DIV2K dataset with $\times$4 super-resolution. On the right, we can observe the maps produced by our perceptual model.
@@ 2 Related Work
Single-frame super-resolution Super-Resolution (SR) or image-upsampling has traditionally been addressed through interpolation-based techniques. Nearest-Neighbor, Bilinear or Bicubic interpolation techniques are commonly used, featuring increasingly bigger receptive fields and perceived quality, at the cost of performance. However, interpolation-based techniques that solely consider the input image are fundamentally limited in terms of upsampling capacity, as high frequency signals lost in the downsampled or compressed image could never be recovered from the signal itself. Leveraging natural image statistics, either explicitly [44] or implicitly [39] has shown greater potential, as lost frequencies can be composed from the expected distribution already present in natural images. Implicit (Deep-Learning based) methods have recently received the greatest attention, as large datasets of high resolution images [10, 34] and advances in neural-network architectures [19, 14, 12].
@@
Temporally-consistent Video SR Temporally-consistent video SR has received ample attention in the literature, for both pre-rendered video [20] and real-time content (i.e. videogames) [30, 15, 2]. The main difference between single image SR is the availability of additional frames, as well as extra information from the rendering engine in the case of videogames (i.e. motion vectors, material information). Traditionally, single-image learning-based approaches struggle to provide consistent SR across frames due to the implicit, difficult-to-interpret latent space they leverage, which does not guarantee temporally consistent outputs. To ensure smoothness across frames, most works simply condition the current frame upsampling on previous or subsequent frames through motion vectors or optical flow to ensure smoothness [30].
There is a fundamental difference between the traditional approach to video SR and our proposal that is important to clarify. Traditionally, video SR methods aim to achieve a better spatial reconstruction due to the availability of multiple frames. However, human sensitivity to spatial details decreases with movement, so a human vision centric approach should leverage this and reduce quality as a funtion of movement magnitude, without noticeable visual degradation. Our proposed model quantifies the loss of visual acuity with motion, and appropriately decreases the spatial quality of SR. Therefore, we can get faster per-frame SR with videos, in a manner adaptive to factors such as the nature of the content and amount of movement.
@@ 3.1 Low-Level Human Vision
Owing to evolutionary imperatives of efficiency, the human visual system has evolved to be a compressive system. What this essentially means is that we are not able to resolve all the visual information that is coming into our eyes. The first critical aspect is that due to center surround receptive field of the early visual system, we are much more sensitive to variations in contrast rather than absolute luminance [35]. Furthermore, it is well known that human contrast perception is highly dependent on factors such as spatial frequency and luminance. Due to the nature of the collective receptive field our early human vision, we are most sensitive to a narrow band of spatial frequencies, and our sensitivity falls of at lower and higher spatial frequencies. This phenomena is aptly captured by a model of the early visual system called the Contrast Sensitivity Function (CSF) [5]. The CSF is highly dependent on factors such as local adapting luminance, size of the stimulus, eccentricity, and the amount of movement. For example, universal sensory models such as the Weber-Fechner law tell us that our ability to detect contrast decreases with increasing luminance [26]. Furthermore, due to effects such as contrast masking, neighboring contrast is known to strongly effect human visual perception [37]. This is why we are sometimes less likely to see a loss of resolution in highly textured areas as opposed to independent strong edges. Another very important aspect is that movement or temporal variation decreases our ability to detect and resolve contrast [3]. This behavior is quantified by the dependency of the CSF on temporal frequencies as well as spatial frequencies.
@@ 3.2 Visual Difference Predictors (VDP)
Inspired by the compressive nature of human vision, Visual Difference Predictors (VDP) are models that aim to predict the perceived differences between two images based on robustly modeling the frequency selective nature of the early visual system. The framework was originally introduced by Daly et al. [9]. Since then, there have a lot of improvements and extensions to the original model. [26] designed the HDR-VDP, which was one of the first human perception inspired metrics aim to quantify visual differences between HDR images. [27] then extended the VDP to account for human peripheral vision and color [29]. [38] employed the VDP framework to control spatial quality in VR-HMDs. [37] designed a variation of the VDP framework for real-time perceptually optimized tone mapping. Our perceptual model can be thought of as a VDP framework specifically tailored for real-time application to the problem of Deep Learning based Super-Resolution. To the best of our knowledge, our framework is the first application of robust perceptual frameworks to efficient neural network based image/video processing.
4 Method
Our approach is centered around the frequency domain interpretation of super-resolution and the well-established fact that the early visual system is frequency-selective. The main motivation is that the difference between low and high-resolution images lies in the attenuation and removal of higher spatial frequencies. The task of an SR neural network is to reconstruct the missing high spatial frequencies. The better and stronger the neural network, the better the reconstruction. The main idea behind our work is that due to the limitations of the human visual system, we do not always need a perfect reconstruction. If we can quantify the spectral nature of an SR method, we can guide the method using models of visual perception to deliver the least expensive reconstruction required for optimal visual quality, i.e., any further improvement in reconstruction leads to no perceptual benefit while wasting computational resources.
@@ 4 Method
Given different variants of an SR method, our method aims to predict which version should be used in a specific region of an image or video frame. We propose to use attenuation curves to first quantify the reconstruction capability of a given variant. The attenuation curves express the ratio of the radially averaged 2D Fourier transform of the reconstructed output and its full-resolution counterpart. We demonstrated that such curves can be computed on a set of images and reused. Furthermore, we design a perceptual model that expresses the required reconstruction quality in the form of the attenuation curve. Later, our method selects an appropriate SR variant to ensure adequate reconstruction quality for a region while minimizing computational costs. In practice, our method works on image patches which are both input to the SR method and our prediction. Below, we describe all the components of the method followed by the results.
@@ 4.1 Attenuation Response Estimation
For a given SR method and an input image, we can quantify the quality of reconstruction by comparing the magnitude of the Fourier Transform of the reconstructed image and its ground-truth version. Given a ground-truth image $I$, we first downscale it by a factor $k$, producing image $I{\downarrow_{k}}$. Later, we use a super resolution method to upscale that image back to its original resolution. Comparing the Fourier transform of the resulting image $\hat{I}$ and the ground-truth counterpart $I$, allows us to quantify the reconstruction power of the analyzed SR method. More formally, given a super-resolution method $\phi$ and a test image $I$, we define the frequency dependent attenuation curve as:
$\alpha^{\phi}_{k}(I,f)=\frac{|\mathscr{F}\left(\phi(I{\downarrow_{k}})\right)(f)|}{|\mathscr{F}\left(I\right)(f)|},$
@@
Although not guaranteed, the value of $\alpha_{k}^{\phi}$ is expected to lie on $(0,1)$ range, where $\alpha_{k}^{\phi}(f)=0$ means that the SR method was not able to reconstruct content at spatial frequency $f$, while $\alpha_{k}^{\phi}(f)=1$ indicates the full capability in reconstructing this part of the signal. It has to be noted that this measure quantifies the presence of the signal in the reconstructed output and not its correctness. Nevertheless, we use this as a proxy for the reconstruction quality. This choice was further motivated by the fact that, although neural networks are non-linear functions, super-resolution is a low-level task with consistent and deterministic characteristics in the frequency domain. The average spectrum of natural images is known to adhere to consistent Fourier characteristics like the inverse power law fall-off. Thus, an average attenuation curve is a good approximation for network response to a general natural image for the task of SR. In order to be more conservative, there is a possibility to use a particular percentile or quartile around the average, but our empirical analysis did not reveal it necessary, which is later confirmed by the results of our user experiments.
@@
For all the variants of the super-resolution techniques considered in this work, we pre-compute the attenuation response curves using the above procedure. We always use 19 natural images from the set5 [6] and set14 [46] datasets, and compute different sets of curves for downscaling parameter $k\in\{2,4,8\}$. Fig. 2 presents a set of curves for the case of varying the network depth. After estimating the attenuation curves for variants of SR methods, the next step is to estimate which of them is ideal for a given patch of the image or video frame.
@@
Figure 2: The attenuation curves derived for the case of modulating the depth (number of layers) of the network. Depth equal 0 corresponds to a bicubic upsampling. As the depth of the network is reduced the performance of the solution in reconstructing high-spatial-frequency signal is reduced. The units of the spatial frequency in the plot are cycles-per-pixel.
4.2 Perceptual Contrast Modeling
The first step is to convert normalized pixel values into physical luminance units as seen by the visual system. We achieve this by first linearizing the input intensity values, and applying a forward display model [25] as:
$L_{disp}(p)=(I_{in}(p))^{\gamma}\cdot(L_{max}-L_{black})+L_{black}+L_{r}$ (1)
where $L_{disp}(p)$ is the displayed luminance in nits, $I_{in}(p)$ is the normalized pixel intensity value for the pixel $p$, $\gamma$ is the display gamma, $L_{max}$ is the peak brightness of the display, $L_{black}$ is the black level of the display, and $L_{r}$ is the ambient light reflected from the display surface.
Our visual system is attuned to detect and process contrast, and the efficacy depends on factors such as spatial frequency, content viewing distance, adapting luminance, and amount of motion. Models such as the CSF (which are the base for our VDP framework) usually study the perception of Michelson contrast of Gabor patches with different characteristics [27] . The Michelson contrast of a Gabor patch can be defined as:
$C_{michelson}=\frac{L_{max}-L_{min}}{L_{max}+L_{min}}$ (2)
In order to get a very fast multi-scale representation of pixel-wise Michelson contrast in the input image, we construct a contrast pyramid [27]. A contrast pyramid level is an element-wise ratio between a Laplacian pyramid ($\Delta L(f,p)$) and Gaussian pyramid level ($L_{a}(f,p)$). Each level is centered around a band of spatial frequencies, and encodes contrast in the specific range:
$C(f,p)=\frac{\Delta L(f,p)}{L_{a}(f,p)+\epsilon}$ (3)
and
$f=2^{-(l+0.5)}\cdot ppd$ (4)
where $f$ is the spatial frequency in cycles per visual degree, and $p$ is the pixel index, $l$ is the pyramid level number, and $ppd$ is the pixels per visual degree that the image spans on the human eye (dependent on display dimensions and viewing distance). It is important to point out that the Gaussian pyramid level used is one higher than the Laplacian pyramid, to account for the local luminance adaption region size in human vision (approx 0.5-1 visual degrees) [37]. The contrast pyramid encodes contrast in physical units, the next step is to perceptually normalize it. As we are interested in the detectability of changes in contrast, our problem is more near-threshold than supra-threshold. For near threshold contrast perception modeling, we can normalize physical contrast by multiplication with the corresponding CSF value.
$C_{n}(f,p)=C(f,p)S_{CSF}\left(f,\omega,L_{a}(f,p)\right)$ (5)
where $\omega$ is the temporal frequency in Hertz. Now, as the CSF is usually measured with Gabor patches on gray backgrounds, it does not account for the visual masking neighboring contrast does to reduce visibility. We model contrast masking as visibility threshold elevation due to the neighboring contrast through a transducer function. The important aspect to remember is that the strongest masking signal is the octave closest to the base contrast, so we only need to consider masking in the same contrast pyramid level as the base contrast [45]. So the final perceptual contrast measure in Just Noticeable Difference (JND) units comes out to be:

### v2

# Human Vision Constrained Super-Resolution

@@ Abstract
Abstract
This supplementary file presents further details and additional results of the proposed method. These results illustrate further outcomes of the user study. Subsequently, the preliminary prototype of the method applied to AR/VR headsets are illustrated. Finally, further qualitative results of the perceptual model are presented to demonstrate the efficacy of the method.
@@ 1 Introduction
Super-resolution (SR) has quickly become a fundamental tool in imaging and media distribution, given the increasing requirements of delivering higher quality content at lower bandwidths, and as general compression tools to deal with escalating imaging sensor resolutions. In media production such as virtual reality, augmented reality, and video games, SR is indispensable to cater to the high-resolution, high-framerate requirements of modern displays and low power budgets [39, 36, 2, 18]. With the advent of hardware-specific accelerators for efficiently running DL models [37], most modern approaches, even for real-time needs, involve using convolutional neural networks trained on large priors of natural images, which during run-time take low-resolution images as input and produce higher resolution versions. However, even with ever-increasing computational power at our disposal, the computational burden of high-quality SR methods is still problematic [1]. In fact, state-of-the-art methods on real-time SR using neural networks, such as Bicubic++ [8], directly compete with efficient, classic bicubic interpolation techniques, yielding marginal quantitative improvements while still struggling to compete in runtime. At the same time, the human visual system (HVS) is compressive by nature [59], meaning it has limited capabilities to resolve detail beyond some thresholds determined by viewing conditions, spatial frequencies, color or motion; any further improvement in reconstruction quality achieved by using a more expensive model can be seen as wasted resources. Our key insight is that, given that SR images are to be observed by a human, we can take advantage of these naturally compressive capabilities of the HVS [59] to process differently areas of the input image depending on its characteristics and human sensitivity to those, such as spatial frequency, luminance, color, contrast, or even, in the case of videos, motion. Computational resources are allocated to perceptually meaningful areas, as determined by our low-level visual model; analogously to how lossy compression schemes such as JPEG similarly allocate memory by leveraging the HVS.
@@
This observation is the foundation of our work. First, we quantify the frequency reconstruction capability of baseline SR neural networks after each layer, doing so through attenuation curves. Then, after dividing an image or individual frame into a set of patches, our Human Visual Processing Framework (HVPF) (grounded on recent contrast-sensitivity functions [32]) predicts how many layers of the network should the patch be processed through; visually sensitive patches will be processed by the whole model, whereas the least visually meaningful content will simply resort to bicubic interpolation, Fig. 1 shows a visual example. We apply our model to other use cases (with the goal of computational efficiency) such as selecting a network of appropriate depth from a set of candidates. In all of these cases, the goal is make the inference as fast as possible without noticeable quality degradation. Through a series of user studies and quantitative measurements of quality and runtime, we demonstrate that our HVPF enables faster runtime woth no perceivable loss of quality when compared with the baseline models. Furthermore, contemporary SR methods often only consider foveal vision, making them suboptimal for wide-field-of view systems such as AR/VR displays. We also propose an eccentricity-aware extension of our model for VR and AR applications, where computational efforts are further directed as per the loss of human visual acuity in peripheral vision.  [39].
@@ 2 Related Work
Single-frame SR
Super-Resolution or image-upsampling has traditionally been addressed through interpolation-based techniques. Nearest-Neighbor, Bilinear or Bicubic interpolation techniques are commonly used, featuring increasingly bigger receptive fields and perceived quality, at the cost of performance. However, interpolation-based techniques that solely consider the input image are fundamentally limited in terms of upsampling capacity, as high frequency signals lost in the downsampled or compressed image could never be recovered from the signal itself. Leveraging natural image statistics, either explicitly [54] or implicitly [47] has shown greater potential, as lost frequencies can be composed from the expected distribution already present in natural images. Implicit (Deep-Learning based) methods have recently received the greatest attention, as large datasets of high resolution images [12, 42] and advances in neural-network architectures [24, 17, 14].
@@
Temporally-consistent Video SR
Temporally-consistent video SR has received ample attention in the literature, for both pre-rendered video [25] and real-time content (i.e. videogames) [36, 18, 2]. The main difference between single image SR is the availability of additional frames, as well as extra information from the rendering engine in the case of videogames (i.e. motion vectors, material information). Traditionally, single-image learning-based approaches struggle to provide consistent SR across frames due to the implicit, difficult-to-interpret latent space they leverage, which does not guarantee temporally consistent outputs. To ensure smoothness across frames, most works simply condition the current frame upsampling on previous or subsequent frames through motion vectors or optical flow to ensure smoothness [36].
There is a fundamental difference between the traditional approach to video SR and our proposal that is important to clarify. Traditionally, video SR methods aim to achieve a better spatial reconstruction due to the availability of multiple frames. However, human sensitivity to spatial details decreases with movement, so a human vision centric approach should leverage this and reduce quality as a function of movement magnitude, without noticeable visual degradation. Our proposed model quantifies the loss of visual acuity with motion, and appropriately decreases the spatial quality of SR. Therefore, we can get faster per-frame SR with videos, in a manner adaptive to factors such as the nature of the content and amount of movement.
@@ 3.1 Low-Level Human Vision
Owing to evolutionary imperatives of efficiency, the HVS has evolved to be a compressive system [59]. What this essentially means is that we are not able to resolve all the visual information that is coming into our eyes. The first critical aspect is that due to center surround receptive field of the early visual system, we are much more sensitive to variations in contrast rather than absolute luminance [43]. Furthermore, it is well known that human contrast perception is highly dependent on factors such as spatial frequency and luminance. Due to the nature of the collective receptive field our early human vision, we are most sensitive to a narrow band of spatial frequencies, and our sensitivity falls of at lower and higher spatial frequencies. This phenomena is aptly captured by a model of the early visual system called the Contrast Sensitivity Function (CSF) [6]. The CSF is highly dependent on factors such as local adapting luminance, size of the stimulus, eccentricity, and the amount of movement. For example, universal sensory models such as the Weber-Fechner law tell us that our ability to detect contrast decreases with increasing luminance [30]. Furthermore, due to effects such as contrast masking, neighboring contrast is known to strongly effect human visual perception [45]. This is why we are sometimes less likely to see a loss of resolution in highly textured areas as opposed to independent strong edges. Another very important aspect is that movement or temporal variation decreases our ability to detect and resolve contrast [4]. This behavior is quantified by the dependency of the CSF on temporal frequencies as well as spatial frequencies.
@@ 3.2 Visual Difference Predictors (VDP)
Inspired by the compressive nature of the human vision, Visual Difference Predictors (VDP) are models that aim to predict the perceived differences between two images based on robustly modeling the frequency selective nature of the early visual system. The framework was originally introduced by Daly et al. [11]. Since then, there have a lot of improvements and extensions to the original model. [30] designed the HDR-VDP, which was one of the first human perception inspired metrics aim to quantify visual differences between HDR images. [31] then extended the VDP to account for human peripheral vision and color [33]. [46] employed the VDP framework to control spatial quality in VR-HMDs. [45] designed a variation of the VDP framework for real-time perceptually optimized tone mapping. Our HVPF can be thought of as a VDP framework specifically tailored for real-time application to the problem of Deep Learning based SR. To the best of our knowledge, our framework is the first application of robust Human Vision frameworks to efficient neural network based image/video processing.
4 Our Method: A Human Visual Processing Framework (HVPF)
Our approach is centered around the frequency domain interpretation of SR and the well-established fact that the early visual system is frequency-selective. The main motivation is that the difference between low and high-resolution images lies in the attenuation and removal of higher spatial frequencies. The task of an SR neural network is to reconstruct the missing high spatial frequencies. The better and stronger the neural network, the better the reconstruction. The main idea behind our work is that due to the limitations of the HVS, we do not always need a perfect reconstruction. If we can quantify the spectral nature of an SR method, we can guide the method using models of human visual perception to deliver the least expensive reconstruction required for optimal visual quality, i.e., any further improvement in reconstruction leads to no perceivable benefit while wasting computational resources.
@@ 4 Our Method: A Human Visual Processing Framework (HVPF)
Given different variants of an SR method, our method aims to predict which version should be used in a specific region of an image or video frame. We propose to use attenuation curves to first quantify the reconstruction capability of a given variant. The attenuation curves express the ratio of the radially averaged 2D Fourier transform of the reconstructed output and its full-resolution counterpart. We demonstrated that such curves can be computed on a set of images and reused. Furthermore, we design a framework that expresses the required reconstruction quality in the form of the attenuation curve. Later, our method selects an appropriate SR variant to ensure adequate reconstruction quality for a region while minimizing computational costs. In practice, our method works on image patches which are both input to the SR method and our prediction. Below, we describe all the components of the method.
@@ 4.1 Attenuation Response Estimation
For a given SR method and an input image, we can quantify the quality of reconstruction by comparing the magnitude of the Fourier Transform of the reconstructed image and its ground-truth version. Given a ground-truth image $I$, we first downscale it by a factor $k$, producing image $I{\downarrow_{k}}$. Later, we use a SR method to upscale that image back to its original resolution. Comparing the Fourier transform of the resulting image $\hat{I}$ and the ground-truth counterpart $I$, allows us to quantify the reconstruction power of the analyzed SR method. More formally, given a SR method $\phi$ and a test image $I$, we define the frequency dependent attenuation curve as:
$\alpha^{\phi}_{k}(I,f)=\frac{|\mathscr{F}\left(\phi(I{\downarrow_{k}})\right)(f)|}{|\mathscr{F}\left(I\right)(f)|},$ (1)
@@
Although not guaranteed, the value of $\alpha_{k}^{\phi}$ is expected to lie on $(0,1)$ range, where $\alpha_{k}^{\phi}(f)=0$ means that the SR method was not able to reconstruct content at spatial frequency $f$, while $\alpha_{k}^{\phi}(f)=1$ indicates the full capability in reconstructing this part of the signal. It has to be noted that this measure quantifies the presence of the signal in the reconstructed output and not its correctness. Nevertheless, we use this as a proxy for the reconstruction quality. This choice was further motivated by the fact that, although neural networks are non-linear functions, SR is a low-level task with consistent and deterministic characteristics in the frequency domain. The average spectrum of natural images is known to adhere to consistent Fourier characteristics like the inverse power law fall-off. Thus, an average attenuation curve is a good approximation for network response to a general natural image for the task of SR. In order to be more conservative, there is a possibility to use a particular percentile or quartile around the average, but our empirical analysis did not reveal it necessary, which is later confirmed by the results of our user experiments.
@@
For all the variants of the SR techniques considered in this work, we pre-compute the attenuation response curves using the above procedure. We always use 19 natural images from the set5 [7] and set14 [56] datasets, and compute different sets of curves for downscaling parameter $k\in\{2,4,8\}$. Fig. 2 presents a set of curves for the case of varying the network depth. To obtain a more compact representations of the attenution functions, we model them using Gaussian fall-off:
$\alpha^{\prime}(f)=\frac{1}{a\sqrt{2\pi}}\exp\left(-\frac{(f-b)^{2}}{2a^{2}}\right)+c,$ (3)
where $f$ is the spatial frequency and $a,b,c$ are parameters estimated via fitting. After estimating the attenuation curves for variants of SR methods, the next step is to estimate which of them is ideal for a given patch of the image or video frame.
@@
Figure 2: The attenuation curves derived for the case of modulating the depth (number of layers) of the network. Depth equal 0 corresponds to a bicubic upsampling. As the depth of the network is reduced the performance of the solution in reconstructing high-spatial-frequency signal is reduced.
4.2 Perceived Contrast Modeling
We model the perceived luminance contrast following the VDP framework e.g. [46]. Please refer to the original work for more details on perceptual difference modeling. To summarize our model designed for application to SR, luminance contrast is computed as $C(f,p)$ using a multi-scale Laplacian-Gaussian pyramid, where $p$ is the location and $f$ is spatial frequency [40, 9]. The contrast measure is then normalized by the contrast sensitivity function (CSF), yielding $C_{n}(f,p)$. Finally, the perceived contrast, $C_{t}(f,p)$, also incorporates the visual masking model [55] with parameters $\alpha=0.7$ and $\beta=0.2$.
4.3 Optimization
Given an input patch, the goal is to find the maximum attenuation that is under the resolution capabilities of human vision. According to the contrast model, the attenuation remains undetectable by an observer if the contrast difference between the original image and the attenuated one is under 1 JND. Consequently, to find the attenuation that results in maximum performance gains yet imperceptible quality loss, we optimize the attenuation curve such that it results in exactly 1 JND difference, i.e.,
$\forall_{f}\,\,C_{t}^{\prime}(p,f)-C_{t}(p,f)=1,$ (4)
where $C_{t}(p,f)$ represents the perceived contrast of the input image patch and $C_{t}^{\prime}(p,f)$ represents the perceived contrast of the network output. Assuming the attenuation curves are a response of the network/branch to the input, the attenuation is a modulation of the physical image contrast at different frequencies:
$\alpha^{\prime}(f)=\frac{C^{\prime}(f,p)}{C(f,p)}=\frac{C^{\prime}_{n}(f,p)}{C_{n}(f,p).}$ (5)
Our immediate goal is to estimate the tolerable output contrast $C^{\prime}_{n}(f,p)$ from the constraint in Eq. 4. We start by developing Eq. 4 by substituting expressions from [46], which leads to the following form of the constraint: